The Drama Club hosted a 2003 minute long dance marathon. It started at 7:00 pm Friday. On Sunday, at what time did it end?
A.4:23 am B. 4:38 am C. 6:23 am D. 6:30 am E. 6:38 am

Answer by richwmiller(17219)

2003/60=33 hrs 23 minutes
33-24=9
but Midnight Friday -7:00 pm =5 hrs
9-5=4
4:23 am Sunday